# Linear response and exact hydrodynamic projections in Lindblad equations with decoupled Bogoliubov hierarchies

### v1

@@ I Introduction
Understanding the effects of dissipation on the dynamics of many-particle systems is an important challenge. Significant simplifications occur in cases where the environmental degrees of freedom can be treated in a Markovian approximation, which allows a description in terms of a Lindblad equation (LE)[1, 2, 3]. Many-particle LEs are generally difficult to analyze and most studies are based on perturbative [4, 5] or numerical approaches [6, 7, 8, 9, 10]. This poses the question whether there are examples in which exact results can be obtained. One such class of models are LEs that can be written as imaginary-time Schrödinger equations with non-Hermitian “Hamiltonians” that are quadratic in fermionic or bosonic field operators [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. These are very special in that the density matrix retains its Gaussian form under time evolution. A wider and richer class of exactly solvable LEs was identified more recently with the discovery of Lindbladians that can be mapped to interacting Yang-Baxter integrable models [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Some models were found to exhibit operator-space fragmentation [32, 33, 34] – the Lindbladian analog of Hilbert-space fragmentation [35, 36]– which gives rise to novel kinds of Yang-Baxter integrability and free fermion structures respectively. In a separate development unrelated to integrability LEs that give rise to decoupled BBGKY hierarchies were identified [37, 38, 39, 40, 41, 42, 43, 44, 45], which hugely simplifies the analysis of correlation functions. Importantly, these models typically cannot be mapped onto LEs with Lindbladians that are quadratic in bosons or fermions, and do not preserve Gaussianity of density matrices under time evolution. Interestingly some LEs are both Yang-Baxter integrable and exhibit decoupled BBGKY hierarchies [21].
@@
The purpose of our work is to exploit the simplifications afforded by decoupled BBGKY hierarchies to address three issues of current interest. First, we obtain analytic results for exact hydrodynamic projections [46]. We analyze LEs with a continuous U(1) symmetry related to particle number conservation and infinite temperature steady states, which exhibit diffusive behaviour at late times. The presence of a conservation law then implies that certain observables will exhibit hydrodynamic power-law tails at late times, which are related to the existence of diffusive eigenmodes of the Lindbladian. A generally unsolved problem is to determine the projection of a given operator onto these diffusive modes. Generically only symmetry arguments are available to gain some information on such projections, see e.g. [47] for a recent example.
@@
Second, we consider linear response functions on top of Lindbladian time evolution [48]. This is relevant in the context of pump-probe experiments, and in particular the question how dissipation affects the line shapes seen in linear response functions. Analyzing linear response on top of Lindbladian time evolution is generally a difficult problem, but as we show in the following, this becomes feasible in a class of models with decoupled BBGKY hierarchies.
@@
Third, we utlize the fact that some LEs with decoupled BBGKY hierarchies are Yang-Baxter integrable while others are not to investigate what effects integrability has. In contrast to unitary quench dynamics [49] integrability in LEs does not imply the existence of conservation laws. This is immediately obvious from the fact that many integrable LEs have unique steady states that are completely mixed, i.e. correspond to infinite temperature density matrices.
@@
The outline of this manuscript is as follows. In section II we introduce the class of LE studied in our work. Section III shows that by an appropriate vectorization the equations of motion in the Heisenberg picture can be mapped onto imaginary time Schrödinger equations with non-Hermitian Hamiltonians. The latter take the form of spin-1/2 fermions with purely imaginary hopping terms and short-ranged interactions. Section IV focuses on the time evolution of operators quadratic in fermions in U(1) symmetric models and presents a detailed analysis of the corresponding eigenvalue spectrum of the non-Hermitian Hamiltonians. In section V closed-form expressions for the hydrodynamic projections of U(1) invariant operators are derived. These describe the late-time behaviour of such operators. Section VI discusses the time evolution in LEs with decoupled BBGKY hierarchies that break particle-number conservation. Here the decoupling results in a triangular form of the BBGKY hierarchy, and we show how to solve the resulting equations of motion for operators quadratic in fermions. Operators cubic in fermions decay exponentially in time in all LEs studied here, and Section VII reports the corresponding decay rates. In contrast, operators quartic in fermions can again display hydrodynamic tails. Section VIII discusses the spectra of the non-Hermitian Hamiltonians relevant to the late time behaviour, focusing in particular on a Yang-Baxter integrable case in which an analytic understanding is possible. In sections IX and X we apply our results to the study of non-equal time correlation functions in the steady state and linear response functions after quantum quenches respectively. Finally, we summarize our results and discuss lines of future enquiry in Section XI.
@@ II Lindblad equations with decoupled Bogoliubov hierarchies
Assuming the jump operators are Hermitian, $L^{\dagger}_{\alpha}=L_{\alpha}$, or come in adjoint pairs, $L^{\dagger}_{\alpha}=L_{\beta}$, the Lindblad equation takes the simpler form
@@
  • Such models are not free (there is no Wick’s theorem) and an intially Gaussian density matrix does not remain Gaussian under time evolution.
@@
All the models have the infinite temperature (equilibrium) state as their steady state.
III Structure of the equations of motion
@@
×$\qquad\qquad\times(-1)^{\sum_{r=1}^{m}\ell_{s}}.$ (14)
@@
    ${\cal H}=$ $-iJ\sum_{j,\tau}\big{(}c^{\dagger}_{j,\tau}c_{j+1,\tau}+{\rm h.c.}\big{)}-\nu\gamma\sum_{j}S^{z}_{j}S^{z}_{j+1}$
@@
    $+\gamma\sum_{j}(-1)^{j}\big{[}(n_{j}-1)P_{j+1}-P_{j}(n_{j+1}-1)\big{]}\ .$ (19)
    A key feature of this Hamiltonian is that in all terms there are at least as many fermion annihilation operators as there are creation operators.
III.1 Formal solution of the equations of motion
@@
${\cal H}|\Phi_{R,\alpha}\rangle$ $=\lambda_{\alpha}|\Phi_{R,\alpha}\rangle\ ,$
$|\Phi_{R,\alpha}\rangle$ $=\sum_{\vec{j}_{n},\vec{\ell}_{m}}{\Phi}(\vec{j}_{n},\vec{\ell}_{m})\prod_{r=1}^{n}c^{\dagger}_{j_{r},\uparrow}\prod_{s=1}^{m}c^{\dagger}_{\ell_{s},\downarrow}|0\rangle\ ,$ (20)
@@
the decay rate of the operator $\hat{\Psi}_{n,m}$.
III.2 Single-fermion operators
@@
IV Dynamics of operators quadratic in fermions in $U(1)$-symmetric models
We next turn to the dynamics of operators quadratic in fermions. It is convenient to treat Model V separately froms models I-IV, as particle number is not conserved.
IV.1 Position representation for the eigenvalue equation with $n_{\uparrow}=n_{\downarrow}=1$
The position representation of the eigenvalue equation (III.1) reads
@@ IV.1 Position representation for the eigenvalue equation with $n_{\uparrow}=n_{\downarrow}=1$
where for Model I and III
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\sigma=\pm 1}}\left(\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\sigma}\right)$
@@
$+\nu\sum_{\sigma=\pm 1}\delta_{x,y+\sigma}\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}\ ,$ (28)
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=-iJ\displaystyle{\sum_{\sigma=\pm 1}}(\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\sigma})$
$-\gamma\sum_{\sigma=\pm 1}(\delta_{x,y}\delta_{x,x^{\prime}+\sigma}\delta_{y,y^{\prime}+\sigma}+\xi\delta_{x,y+\sigma}\delta_{x,y^{\prime}}\delta_{y,x^{\prime}})$
@@
$\Phi_{\alpha}(L+1,y)=\Phi_{\alpha}(1,y)\ ,\ \Phi_{\alpha}(x,L+1)=\Phi_{\alpha}(x,1)\ .$ (32)
@@
IV.1.1 Eigenvalue spectrum for Model I

### v2

@@ I Introduction
Understanding the effects of dissipation on the dynamics of many-particle systems is an important challenge. Significant simplifications occur in cases where the environmental degrees of freedom can be treated in a Markovian approximation, which allows a description in terms of a Lindblad equation (LE)[33, 46, 8]. Many-particle LEs are generally difficult to analyze and most studies are based on perturbative [44, 63] or numerical approaches [16, 7, 15, 69, 68]. This poses the question whether there are examples in which exact results can be obtained. One such class of models are LEs that can be written as imaginary-time Schrödinger equations with non-Hermitian “Hamiltonians” that are quadratic in fermionic or bosonic field operators [56, 21, 57, 13, 35, 42, 62, 67, 49, 1]. These are very special in that the density matrix retains its Gaussian form under time evolution. A wider and richer class of exactly solvable LEs was identified more recently with the discovery of Lindbladians that can be mapped to interacting Yang-Baxter integrable models [51, 71, 60, 61, 55, 9, 58, 70, 17, 38, 37]. Some models were found to exhibit operator-space fragmentation [29, 59, 45] – the Lindbladian analog of Hilbert-space fragmentation [53, 54]– which gives rise to novel kinds of Yang-Baxter integrability and free fermion structures respectively. In a separate development unrelated to integrability LEs that give rise to decoupled BBGKY hierarchies were identified [34, 24, 22, 74, 12, 11, 52, 32, 43, 19, 48, 5], which hugely simplifies the analysis of correlation functions. Importantly, these models typically cannot be mapped onto LEs with Lindbladians that are quadratic in bosons or fermions, and do not preserve Gaussianity of density matrices under time evolution. Interestingly some LEs are both Yang-Baxter integrable and exhibit decoupled BBGKY hierarchies [51].
The purpose of our work is to exploit the simplifications afforded by decoupled BBGKY hierarchies to address three issues of current interest.
@@
$\bullet$ First, we obtain analytic results for exact hydrodynamic projections [20]. We analyze LEs with a continuous U(1) symmetry related to particle number conservation and infinite temperature steady states, which exhibit diffusive behaviour at late times. The presence of a conservation law then implies that certain observables will exhibit hydrodynamic power-law tails at late times, which are related to the existence of diffusive eigenmodes of the Lindbladian. A generally unsolved problem is to determine the projection of a given operator onto these diffusive modes. Generically only symmetry arguments are available to gain some information on such projections, see e.g. [50] for a recent example. In order to put some of our results into context we briefly review the usual arguments pertaining to diffusive late-time dynamics and its relation to conservation laws. Let us consider a lattice model with U(1) symmetry corresponding to particle number conservation. The particle density $n_{m}$ then fulfils a continuity equation
$\frac{d}{dt}n_{m}(t)=j_{m-1}(t)-j_{m}(t)\ ,$ (1)
where $j_{m}$ is the particle current density at site $m$. Coarse-graining turns this into
$\frac{\partial}{\partial t}\rho(t,x)=-\frac{\partial}{\partial x}J(t,x)\ .$ (2)
Using that the expectation value of the current (in a spatially inhomogeneous state) is related to the gradient of the density
$\langle J(t,x)\rangle=-D\frac{\partial}{\partial x}\langle\rho(t,x)\rangle+\dots$ (3)
then gives a diffusion equation for $\langle\rho(t,x)\rangle$. Its solution can be cast in the form
$\langle\rho(t>0,x)\rangle=\int\frac{dk}{2\pi}e^{-Dk^{2}t}\int_{-\infty}^{\infty}dx^{\prime}e^{ik(x-x^{\prime})}\langle\rho(0,x^{\prime})\rangle\ .$ (4)
This implies that the Heisenberg-picture operator at late times has hydrodynamic tails
$\rho(t\to\infty,x)\sim\frac{A(x)}{\sqrt{t}}+\dots$ (5)
[…]
$\frac{d}{dt}\hat{X}={\cal L}[\hat{X}$] (6)
must have eigenvalues in the momentum representation that behave for small momenta like $\lambda(k)=-Dk^{2}+\dots$. The corresponding eigenoperators are referred to as diffusive eigenmodes of the time evolution operator. Their construction is another difficult problem whose solution is not known. In our work we derive explicit expressions for them.
@@
$\bullet$ Second, we consider linear response functions on top of Lindbladian time evolution [10]. This is relevant in the context of pump-probe experiments, and in particular the question how dissipation affects the line shapes seen in linear response functions. Analyzing linear response on top of Lindbladian time evolution is generally a difficult problem, but as we show in the following, this becomes feasible in a class of models with decoupled BBGKY hierarchies.
@@
$\bullet$ Third, we utilize the fact that some LEs with decoupled BBGKY hierarchies are Yang-Baxter integrable while others are not to investigate what effects integrability has. In contrast to unitary quench dynamics [30] integrability in LEs does not imply the existence of conservation laws. This is immediately obvious from the fact that many integrable LEs have unique steady states that are completely mixed, i.e. correspond to infinite temperature density matrices.
@@
The outline of this manuscript is as follows. In section II we introduce the class of LEs studied in our work. Section IV shows that by an appropriate vectorization the equations of motion in the Heisenberg picture can be mapped onto imaginary time Schrödinger equations with non-Hermitian Hamiltonians. The latter take the form of spin-1/2 fermions with purely imaginary hopping terms and short-ranged interactions. Section V focuses on the time evolution of operators quadratic in fermions in U(1) symmetric models and presents a detailed analysis of the corresponding eigenvalue spectrum of the non-Hermitian Hamiltonians. In section VI closed-form expressions for the hydrodynamic projections of U(1) invariant operators are derived. These describe the late-time behaviour of such operators. Section VII discusses the time evolution in LEs with decoupled BBGKY hierarchies that break particle-number conservation. Here the decoupling results in a triangular form of the BBGKY hierarchy, and we show how to solve the resulting equations of motion for operators quadratic in fermions. Operators cubic in fermions decay exponentially in time in all LEs studied here, and Section VIII reports the corresponding decay rates. In contrast, operators quartic in fermions can again display hydrodynamic tails. Section IX discusses the spectra of the non-Hermitian Hamiltonians relevant to the late time behaviour, focusing in particular on a Yang-Baxter integrable case in which an analytic understanding is possible. In sections X and XI we apply our results to the study of non-equal time correlation functions in the steady state and linear response functions after quantum quenches respectively. Finally, we summarize our results and discuss lines of future enquiry in Section XII.
@@ II Lindblad equations with decoupled Bogoliubov hierarchies
If the set $\{L_{\alpha}\}$ of jump operators is closed under Hermitian conjugation the Lindblad equation takes the simpler form
@@
  • Such models are not free (there is no Wick’s theorem) and an initially Gaussian density matrix does not remain Gaussian under time evolution.
@@
All the models have the infinite temperature (equilibrium) state as their steady state. Models III and IV are chosen in order to give a clear picture of how the spatial structure of the jump operators affects the properties of the hydrodynamic modes, without making the analysis overly tedious.
III Overview of main results
In order make the manuscript easier to read we now provide a brief overview of the main results derived in the remainder of the manuscript.
III.1 Exact expressions for diffusive eigenoperators
Models I-IV have eigenoperators of the dual Lindbladian that exhibit diffusive behaviour at small momentum
${\cal L}^{*}[\hat{\Phi}(p)]=\lambda(p){\hat{\Phi}}(p)\ ,\quad\lambda(p\to 0)=-Dp^{2}+\dots\ .$ (18)
We derive explicit expressions for these eigenoperators in the sector with one fermion creation and annihilation operator each. These have the form
${\hat{\Phi}}(p)=\sum_{j,\ell}e^{ip\frac{j+\ell}{2}}\phi(p;j-\ell)c^{\dagger}_{j}c_{\ell}\ ,$ (19)
where the amplitudes $\phi(p;j-\ell)$ are model-dependent. In all cases they describe ”particle-hole bound states” in the sense that $\phi(p;j-\ell)$ decay exponentially with respect to the distance $|j-\ell|$. The corresponding decay length, or size of the bound state, tends to zero in the zero momentum limit $p\to 0$.
III.2 Exact hydrodynamic projections
The existence of diffusive eigenoperators implies the presence of long-time tails in the expressions of certain Heisenberg-picture operators. We derive explicit expressions for these hydrodynamic projections. In particular we show that
$\big(c^{\dagger}_{x_{0}}c_{y_{0}}\big)(t\gg\gamma^{-1})$ $\simeq\sum_{x,y}\psi_{x_{0},y_{0}}(x,y;t)c^{\dagger}_{x}c_{y}\ ,$ (20)
where the amplitudes $\psi_{x_{0},y_{0}}(x,y;t)$ are model-dependent, but are all expressed in terms on confluent hypergeometric functions, cf. (109). At asymptotically late times they reduce to (decaying) power-laws in $t$, with exponents that depend on $X=|x-y|+|x_{0}-y_{0}|$ and $Y=(x-x_{0})+(y-y_{0})$
$\psi_{x_{0},y_{0}}(x,y;t)\Bigg|_{\rm leading}=\begin{cases}A_{e}\ (Dt)^{-\frac{1+X}{2}}&\text{Y even}\ ,\\ A_{o}\ (Dt)^{-\frac{2+X}{2}}&\text{Y odd}\ .\end{cases}$ (21)
Here $D$ is model-dependent.
III.3 Non-equal time steady-state correlations
The steady state of models I-IV is a simple infinite-temperature state. We show that equilibrium dynamics in this state can be efficiently analyzed. In particular, signatures of the hydrodynamic modes can be seen in dynamical correlation functions. We show this for the case of the Fourier transform of the density-density correlation function
$G(q,\omega)$ $=\sum_{q}\int_{0}^{\infty}\mathrm{d}t\ e^{i(\omega t-qx)}\left[\frac{1}{2^{L}}\mathrm{Tr}\left[n_{j}(t)n_{\ell}\right]-\frac{1}{4}\right]$
$\sim\frac{A}{i\omega+Dq^{2}}\ ,\ \omega,q\to 0.$ (22)
III.4 Linear response functions in non-equilibrium states
Linear response on top of non-equilibrium time evolution describes certain ”pump-probe” experiments. We analyze this setting for Lindblad equations with decoupled BBGKY hierarchies. In particular we consider the case of the density response to an infinitesimal density perturbation, described by the Lindblad equation
$\frac{d\rho(t)}{dt}={\cal L}_{t}[\rho(t)]={\cal L}_{0}[\rho(t)]-i\sum_{\ell}\xi_{\ell}(t)[n_{\ell},\rho(t)]\ .$ (23)
The linear response is given by the term linear in $\xi_{j}(t)$
${\rm Tr}[n_{j}\rho(t)]-{\rm Tr}[n_{j}\rho_{0}(t)]=\int_{0}^{\infty}dt^{\prime}\sum_{\ell}\xi_{\ell}(t^{\prime})\chi(j,\ell;t,t^{\prime}),$ (24)
where $\rho_{0}(t)$ in density matrix in absence of the perturbation. We determine the Fourier transform of the susceptibility $\chi(j,\ell;t,t^{\prime})$ for initial density matrices corresponding to the ground state of dimerized tight-binding models and show how dissipation affects the dynamical response.
IV Structure of the equations of motion
@@
$\qquad\qquad\times(-1)^{\sum_{s=1}^{m}\ell_{s}}.$ (27)
@@
    ${\cal H}=$ $-iJ\sum_{j,\sigma}\big(c^{\dagger}_{j,\sigma}c_{j+1,\sigma}+{\rm h.c.}\big)-\nu\gamma\sum_{j}S^{z}_{j}S^{z}_{j+1}$
@@
    $+\gamma\sum_{j}(-1)^{j}\big[(n_{j}-1)P_{j+1}-P_{j}(n_{j+1}-1)\big]\ ,$ (32)
    where we have defined $n_{j}=n_{j,\uparrow}+n_{j,\downarrow}$. A key feature of this Hamiltonian is that in all terms there are at least as many fermion annihilation operators as there are creation operators.
IV.1 Formal solution of the equations of motion
@@
the decay rate of the operator $\hat{\Psi}_{n,m}$. The structure of the right eigenvalue equation in models I-IV is simple because
  1. 1.
    The vacuum state $|0\rangle$ defined by $c_{j,\sigma}|0\rangle=0$ is an exact eigenstate with eigenvalue zero;
  2. 2.
    The Hamiltonians commute with the number operators of spin up and spin down fermions
    ${\cal H},N_{\sigma}]=0\ ,\quad N_{\sigma}=\sum_{j}n_{j,\sigma}\ .$ (37)
    This means we can consider eigenstates in sectors with fixed numbers $N_{\sigma}$ of fermions with spin $\sigma$, i.e.
    $|\Phi_{R,\alpha}\rangle=\sum_{\vec{j}_{N_{\uparrow}},\vec{\ell}_{N_{\downarrow}}}{\Phi}_{\alpha}(\vec{j}_{N_{\uparrow}},\vec{\ell}_{N_{\downarrow}})\prod_{r=1}^{N_{\uparrow}}c^{\dagger}_{j_{r},\uparrow}\prod_{s=1}^{N_{\downarrow}}c^{\dagger}_{\ell_{s},\downarrow}|0\rangle.$ (38)
In model V the fermion vacuum $|0\rangle$ is still an exact eigenstate, and while particle number is no longer conserved the structure of the Hamiltonian implies that eigenstates have the form
[…]
where in the sums over $n$ and $m$ the difference $n-m=N_{\uparrow}-N_{\downarrow}$ is kept fixed.
IV.2 Single-fermion operators
@@
V Dynamics of operators quadratic in fermions in $U(1)$-symmetric models
We next turn to the dynamics of operators quadratic in fermions. It is convenient to treat Model V separately from Models I-IV, as particle number is not conserved. We recapitulate the steps set out above:
  1. 1.
    We consider operators of the form (25) with $n+m=2$, e.g.
    $\hat{\Psi}_{1,1}=\sum_{j,\ell}\Psi(j,\ell)c^{\dagger}_{j}c_{\ell}\ .$ (43)
  2. 2.
    The vectorized form of the equation of motion in the Heisenberg picture then becomes, cf. (27)
    $\frac{d}{dt}|\Psi_{1,1}(t)\rangle={\cal H}|\Psi_{1,1}(t)\rangle\ .$ (44)
    This is a 2-particle imaginary time Schrödinger equation with non-Hermitian Hamiltonian.
  3. 3.
[…]
    ${\cal H}|\Phi_{R,\alpha}\rangle=\lambda_{\alpha}|\Phi_{R,\alpha}\rangle.$ (45)
    These are most easily constructed in the position representation, which we discuss in detail in the following. Some of the corresponding steps in the momentum representation are briefly summarized in Appendix A.
V.1 Position representation for the eigenvalue equation with $N_{\uparrow}=N_{\downarrow}=1$
The position representation of the eigenvalue equation (33) reads
@@ V.1 Position representation for the eigenvalue equation with $N_{\uparrow}=N_{\downarrow}=1$
where for Models I and III
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=$ $-iJ\displaystyle{\sum_{\tau=\pm 1}}\left(\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\tau}\right)$
@@
$+\nu\sum_{\tau=\pm 1}\delta_{x,y+\tau}\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}\ ,$ (47)
@@
${\mathfrak{H}}(x,y|x^{\prime},y^{\prime})=-iJ\displaystyle{\sum_{\tau=\pm 1}}(\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}}+\delta_{x,x^{\prime}}\delta_{y,y^{\prime}+\tau})$
$-\gamma\sum_{\tau=\pm 1}(\delta_{x,y}\delta_{x,x^{\prime}+\tau}\delta_{y,y^{\prime}+\tau}+\xi\delta_{x,y+\tau}\delta_{x,y^{\prime}}\delta_{y,x^{\prime}})$
@@
$\Phi_{R,\alpha}(L+1,y)=\Phi_{R,\alpha}(1,y)\ ,\ \Phi_{R,\alpha}(x,L+1)=\Phi_{R,\alpha}(x,1)\ .$ (51)
@@
V.1.1 Eigenvalue spectrum for Model I